OpenStudy (anonymous):

Simplify the expression. x^14/x^7

OpenStudy (solomonzelman):

when you divide you subtract the exponents. \[x^{14}~/~x^7~~=~~x^{14-7}~=~?\]

OpenStudy (solomonzelman):

Do you get this?

OpenStudy (anonymous):

x^7?

OpenStudy (solomonzelman):

YES, BINGO!!!!!!!!!!!!!

OpenStudy (anonymous):

Thanks(: I've got so many more but I don't wanna be a pest...but is it ok if u help with a few more?

OpenStudy (solomonzelman):

\[\huge\color{blue}{a^{-b}=\frac{1}{a^b}}\](if 'b" is a positive number)\[\huge\color{blue}{x^a~/~x^b~=~x^{a-b}}\]\[\huge\color{blue}{x^a~ \times ~x^b=x^{a+b}}\]